2.-Calculate the number of moles of CH4 in a 5.00 L cylinder if the temperature is 26.0°C, and pressure is 790. mmHg.
Answer #1

We use the ideal ags equation to get: number of moles (n) = PV / RT, where P is pressure in atm (790 / 760 = 1.04 atm), V is volume in liters, T is temperature in K ( 273 + 26 = 299 K) and R is gas constant

n = 1.04 * 5 / (0.0821 * 299) = 0.212 mol